OpenStudy (klimenkov):

How many non-negative integer solutions does this equation have? \[x_1+x_2+\ldots+x_N=n\]

OpenStudy (cruffo):

Are you considering the solutions to be ordered or unordered? For example, if we wee trying to solve: x + y + z + w = 12 Would the solution (x,y,z,w) = (9,1,1,1) = (1,9,1,1) = (1,1,9,1) = (1,1,1,9) or would that be considered 4 separate solutions? If they are considered the same, then this is a partitions problem, if they are different, then this is a combinations problem.

OpenStudy (klimenkov):

They are different.

OpenStudy (experimentx):

woops!! that's for ... a+b+c+d = N

OpenStudy (cruffo):

For the smaller example: x + y + z + w = 12 The number of solutions can be determined by viewing the problem in the following way: If 12 units (denoted as 12 u's) seperated by 11 spaces (denoted by 11 s's) are lined up, usususususususususususu and we choose any 3 of the s's and let the others disappear, for example: u u s u u u usu u u u usu then the remaining s's seperate the units into four batches. The number of units in these batches can be used as the values of x,y,z, and w.

OpenStudy (experimentx):

a+b = n would make n+1 solutions a +b + c = n would make